byakuyatogami1234:

Plz help me solve this inequality step by step. 17x+5≤ 70

Extrinix:

Ok so, first off let's start by thinking of it as an equation, not an inequality You must get x by itself \(17x+5 \le 70\) Now to start off we subtract 5 from each side \(17x \le 65\) Now let's divide both sides by 17 \(x \le \hspace{5px}\approx 3.82\) Now, because we divided, we have to switch the sign, making it this \(x \ge 3.82\)

byakuyatogami1234:

@extrinix wrote:
Ok so, first off let's start by thinking of it as an equation, not an inequality You must get x by itself \(17x+5 \le 70\) Now to start off we subtract 5 from each side \(17x \le 65\) Now let's divide both sides by 17 \(x \le \hspace{5px}\approx 3.82\) Now, because we divided, we have to switch the sign, making it this \(x \ge 3.82\)
tysm!!!!!!!!!!

Extrinix:

May I correct myself* It would be \(x \le 3.82\) because it was divided by a positive number, not a negative.

byakuyatogami1234:

@extrinix wrote:
May I correct myself* It would be \(x \le 3.82\) because it was divided by a positive number, not a negative.
oh okie thanks

Extrinix:

Non forsit! c: